OpenStudy (anonymous):

Determine the zeros of the function x2 − 25 = 0

OpenStudy (anonymous):

x^2-25-0 x^2=25 x=±5

OpenStudy (the_fizicx99):

Factor it, (x - 5)^2 = (x + 5)(x - 5) x + 5 = 0 x - 5 = 0 Same as \(\ \sf \pm 5 \)

OpenStudy (anonymous):

thank you guys it really help me :)